General Meeting Summary 3/17/22

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Fernando, Julie, Petr, Brad, Paul K., Steven L., David H., Hamlet, Hakob, Rolf, Vladimir, Tanja, Aaron, Arshak, Carlos, Bogdan, Erik W., Marie, 757-269-7191, Vardan, Alexandre,


REVIEW ACTION ITEMS FROM COLLABORATION MEETING

HV DIVIDERS

  • Successful communication with vendor and assembly company
  • After a number of efforts, procurement received confirmation from SAMTEC that the connectors will be shipped on Monday (2/28). Procurement also requested expedited shipping so we can ship these to the assembly company this coming week.
  • sent the assembly company a set of spare connectors purchased for the 1st articles (10 units). Received 10 first articles and they are of good quality. Tests confirm the prototype measurements. The company will be notified to proceed.
  • removed all sockets (1140) from the old PCbs. As 1080 will be installed, we will have over 5% complete spares plus another 6% spare PCBs w/o sockets. Our goal is to have 1080 tested and ready for installation.
  • Expected assembly in April - three people committed to start assembling the sockets once we receive the boards from the assembly company
  • Expect all ready in middle of May


CRYSTALS AND ASSEMBLY PREPARATIONS

  • Reflector production is on track for having all foils preshaped by mid-May
  • Crystal storage space and cabinets/shelves
  • Brad and Vladimir visited the assembly space and identified storage space for wrapped crystals
  • Crystal Wrapping - aim to complete as much as possible in May to be ready when Orsay team arrives
  • DSG available to help with 1-2 people in late May/June
  • Spreadsheet for crystal/module tracking has been developed
  • column headers were collected by NPS collaborators
  • spreadsheet with four tabs to track different aspects (module, crystal properties, electronics, etc) was generated by JLab admin personnel (Jennifer, Sadie)
  • send comments for improvement to hornt@cua who will collect and pass on
  • Initial spreadsheet with onsite personnel developed
  • needs reviewed and completed as needed by NPS collaborators
  • Action Items:
  • email Amrit/Patrizia for official confirmation of DSG participation
  • post spreadsheets (NPS module tracking, onsite personnel)


DAQ/ELECTRONICS AND SOFTWARE

  • D. Hamilton gives on overview of the recent DAQ/Electronics and Software meetings (details on agenda and summaries can be found on the NPS Wiki Meetings page)
  • two meetings so far
  • in the first meeting D. Hamilton agreed to be the software lead. Brad is the overall coordinator. Paul K. and Steve W., et al. offered to help with various tasks
  • good overall plan is in place - what needs to be ready and test plans for detector checkout
  • Important for the overall approach is to develop tools that are as close as possible to those used during production, e.g., ET based display - idea is to test entire hardware and software chain early
  • A new development from the FEG was mentioned that may simplify event blocking
  • follow up on this is needed
  • for now will proceed with the present plan


TRAVEL TO JLAB

  • JLab now in Medcon3
  • For people new to JLab: need all training - most of it can be done online
  • Radworker I training - also need practical
  • Based on previous experience from Orsay team it should take about half a day assuming most of the training has been completed online


DSG UPDATE

  • EPICS Process Variable (PV) list development
  • Ansys thermal simulations - suggests that crystal temperature depends on ambient temperature in the box
  • Developed model of sensor (temperature and humidity) placement in the box
  • ten sensor locations
  • twenty channels to add for temperature and humidity each
  • at the moment have 8 empty channels at the terminal - perhaps not all needed? add additional terminal?
  • High voltage supply cables - glue added to secure connectors
  • also used on Hall A hodoscopes - David can check on what glue was used


MECHANICAL DESIGN/ENGINEERING UPDATE

  • Work on the NPS support structure in the hall - a few photos shown
  • Some corrective action to address difference between drawing and natural deflection of the SHMS
  • Concern about helium release from SHMS quads and possible damage to NPS PMTS
  • Temperature of quads during NPS running needs to be decided still
  • might be able to "park" (80K He) the quads - still large flow of He through magnets though - to reduce any issues for calorimeter concern is quench and emission of He and possible damage to PMTs.
  • Overall, reducing liquid helium flow and not having magnets running should reduce the risk. Only issue remaining is ESR crashing, but normally no major release of He into hall unless more catastrophic power failure
  • Roof block path forward
  • cutting into existing block - cut notch into it (not during this SAD though)


TCS

  • Schedule a dedicated meeting within the next week


NEXT MEETING: THURSDAY 14 APRIL 2022 AT 9:00AM ET